A consulting company estimated market demand and supply in a perfectly competitive industry and obtained the following results:

Qd = 25,000 − 5,000P + 25M

Qs = 240,000 + 5,000P − 2,000PI

where P is price, M is income, and PI is the price of a key input. The forecasts for the next year are ̂ = $15,000 and I = $20.

Average variable cost is estimated to be AVC = 14 − 0.008Q + 0.000002Q2 Total fixed cost will be $6,000 next year. What is the price forecast for next year?

$12

$60

$20

$68
Answer #1

So, here the estimated demand and the supply functions are given. So, at the equilibrium “Qd” must be equal to “Qs”.

=> Qd = Qs, => 25,000 - 5000*P + 25*M = 240,000 + 5,000*P - 2,000*PI, “M=15,000” and “PI=20”.

=> 25,000 - 5000*P + 25*15,000 = 240,000 + 5,000*P - 2,000*20.

=> 400,000 - 5000*P = 200,000 + 5,000*P, => 10,000*P = 200,000, => P = 200,000/10,000.

=> P = $20, => the correct answer is “C”.